# Membership problems in braid groups and Artin groups

### v2

@@ 3.2. Further undecidability in $A(P_{4})$
Let $A(P_{4})$ be the right-angled Artin group (6). Then:
@@
For any finite graph $\Gamma$, [DLS91, Corollary 5] asserts that $A(\Gamma)\leq A(P_{4})$ if and only if $\Gamma$ is a finite forest (reproved in [KK13, Corollary 1.9]). This construction passes by considering the normal closure of $\{b,c\}$ inside $A(P_{4})$, which is isomorphic to a right-angled Artin group on an infinite tree with vertices the conjugates of elements in $A(P_{4})$. In particular, the subgroup of (6) generated by
$aca^{-1},b,c,dbd^{-1}$
is isomorphic to $A(P_{4})$ (with the four vertices of $P_{4}$ in the specified order), and this path extends to a bi-infinite path. Note that in particular $aca^{-1}$ and $c$ are the basis of a free subgroup of rank $2$. On this bi-infinite path, we can find a new copy of $A(P_{4})$, say generated by four elements $v_{1},\dots,v_{4}$, and such that all of the $v_{i}$ lie to the right of $dbd^{-1}$. Furthermore, we can find an element $z$ two steps to the left of $aca^{-1}$, so that $\langle aca^{-1},c,z\rangle$ generates a free group of rank $3$. In particular, we have that $A(P_{4})\ast F_{3}\cong\langle v_{1},\dots,v_{4}\rangle\ast\langle aca^{-1},c,z\rangle$. Thus, we can now apply Lemma 3.4, and both (i) and (ii) follow by taking $x=aca^{-1}$ and $y=c$ in the statement of that lemma, and noting that $xy^{-1}\neq 1$ in $A(P_{4})\ast F_{3}$, and hence $xy^{-1}\in\operatorname{Mon}(\Delta)$ if and only if $xy^{-1}\in\operatorname{Sgp}(\Delta)$. ∎
@@ 3.3. Undecidability in the braid group $\mathbf{B}_{4}$
We end this section by remarking that all the decidability results proved in this section for braid groups also hold true for a closely related object called the pure braid group $\mathbf{PB}_{n}$. Indeed, it well-known and also clear from equation (1) above (and the graphical representation of braids see [KT08, §1.2.2]) that $\mathbf{B}_{n}$ surjects onto $S_{n}$. The kernel of this homomorphism is called the pure braid group $\mathbf{PB}_{n}$, which hence has index $n!$ in $\mathbf{B}_{n}$. All of the undecidability results above for $\mathbf{B}_{n}$, were proved by means of embedding certain right-angled Artin groups $A(\Gamma)$ into $\mathbf{B}_{n}$. However, the results are equally valid for $\mathbf{PB}_{n}$, by the following reasoning. Since all sufficiently high powers of the generators of a right-angled Artin group again generate a right-angled Artin group [CP01], it follows that if $A(\Gamma)\leq\mathbf{B}_{n}$ for some finite graph $\Gamma$, then we also have $A(\Gamma)\leq\mathbf{PB}_{n}$ by considering powers of the generators, where the powers are sufficiently large multiples of $[\mathbf{B}_{n}:\mathbf{PB}_{n}]=n!$. In particular, $A(\Gamma)\leq\mathbf{B}_{n}$ if and only if $A(\Gamma)\leq\mathbf{PB}_{n}$. Thus, all undecidability results above for $\mathbf{B}_{n}$ with $n\geq 4$ also hold for $\mathbf{PB}_{n}$ with $n\geq 4$. To see that all the statements in Theorem 3.6 remain true for the pure braid group $\mathbf{PB}_{4}$, it simply suffices to note that $\gamma_{0}\in\mathbf{PB}_{4}$. Similarly, when $n\leq 3$, since $\mathbf{B}_{n}$ has decidable rational subset membership problem and $\mathbf{PB}_{n}\leq\mathbf{B}_{n}$ it follows that the rational subset membership problem (along with all other decision problems discussion in this section) is decidable in $\mathbf{PB}_{n}$ when $n\leq 3$.
@@ 4. Artin groups
The first result we need to establish is the following.
Theorem 4.1.
Every subgroup separable Artin group has decidable rational subset membership problem.
@@
In [KSS07, Theorem 3.1] it is show that a finitely generated group $G$ has decidable rational subset membership problem if and only if the word problem for $G$ with respect to $A$ is RID (rational intersection decidable). Here the word problem for the group is the set of all words over the finite generating set that are equal to the identity element of the group. In [LS08] Lohrey & Steinberg introduce another class called SLI (semilinear intersection) languages, and corresponding class of SLI-groups. They prove [LS08, Lemma 4] that this class is contained in the class of languages RID, so every SLI-group is an RID-group and hence has decidable rational subset membership problem.
@@
    $A$ has decidable submonoid membership problem;
@@
    $A$ has decidable rational subset membership problem;
@@
((iii) $\Rightarrow$ (ii)) By [AL24, Theorem 1.1] (which is proved using the main result of [AL21]) it follows that if $\Gamma$ does not embed any of the graphs in (a), (b), or (c) then $A(\Gamma)$ is subgroup separable which by Theorem 4.1 implies that $A$ has decidable rational subset membership problem. ((ii) $\Rightarrow$ (i)) is trivial. ((i) $\Rightarrow$ (iii)) Suppose that $\Gamma$ does embed one of the graphs of (a), (b) or (c) as an induced subgraph. Let us consider each case in turn. Case (a): In this case by [CP01] $A(\Gamma)$ embeds the RAAG $A(C_{4})\cong F_{2}\times F_{2}$ and hence contains a fixed finitely generated submonoid (in fact subgroup) in which membership is undecidable by [Mik58]. Cases (b) and (c): In these cases it follows from the argument of the proof of [AL21, Lemma 4.2] that in each case $A(\Gamma)$ embeds [^3] We note that diagram (4.1) in the proof of [AL21, Lemma 4.2] there is an edge between $x$ and $y$ that should not be there. Indeed, by the definition of $R(B)$ in that paper for there to be an edge between $x$ and $y$ one would need every vertex of $\{a,b\}$ to be connected to every vertex of $\{b,c\}$ by a $2$-labelled edge in the defining graph of the Artin group which is not true. This does not affect the validity of the proof of [AL21, Lemma 4.2], or the proof of Theorem 4.4 above, since the RAAG of the graph (4.1) with that edge removed still contains an induced subgraph isomorphic to a path of length $4$. Related to this, note that since the braid group $B_{4}\cong A(2,3,3)$ does not embed $A(C_{4})=F_{2}\times F_{2}$ it follows that (4.1) cannot contain a square as an induced subgraph. the RAAG $A(P_{4})$ and hence by [LS08] it follows that $A(\Gamma)$ contains a fixed finitely generated submonoid in which membership is undecidable. ∎
@@
Let $A=A(\Gamma)$ be an Artin group with undecidable submonoid membership problem, so $\Gamma$ embeds one of the graphs in (a), (b) or (c) from the statement of Theorem 4.4.
@@
    If $\Gamma$ embeds one of the graphs in (b) or (c) then $A(\Gamma)$ embeds the RAAG $A(P_{4})$ and the following problems are all undecidable in $A(\Gamma)$: the rational subset membership problem, submonoid membership problem, fixed-target submonoid membership problem, and the subsemigroup intersection problem.
@@
    If $\Gamma$ embeds one of the graphs in (a) then $A(\Gamma)$ embeds the RAAG $A(C_{4})$ and all the problems listed in (i) are undecidable in $A(\Gamma)$. Additionally, all the following problems are undecidable in $A(\Gamma)$: the group problem, identity problem, and subgroup membership problem.
@@
Each and every one of the following problems is decidable for an Artin group $A=A(\Gamma)$ if and only if $\Gamma$ does not embed one of the graphs in (a), (b) or (c) from the statement of Theorem 4.4: the rational subset membership problem, submonoid membership problem, fixed-target submonoid membership problem, and the subsemigroup intersection problem.
Just as for the braid group $\mathbf{B}_{4}$ we do not know exactly which Artin groups have decidable identity problem or group problem. Similarly, it is unknown exactly which Artin groups have decidable subgroup membership problem. In particular, it remains open whether the braid group $\mathbf{B}_{4}$ has decidable subgroup membership problem. Related to this, it follows from the proof of [AL21, Lemma 4.2] that $\mathbf{B}_{4}$ embeds $A(C_{5})$, where $C_{5}$ is a pentagon, and it is also an open problem whether $A(C_{5})$ has decidable subgroup membership problem; see [LS08, Section 5].

### v3

@@ 3.2. Further undecidability in $A(P_{4})$
To prove the proposition we shall need the following result which is a special case of results proved in [DLS91]. For the convenience of the reader we provide the details on how to deduce this lemma by applying results from [DLS91].
Lemma 3.5.
Let
$G=A(P_{4})=\langle a,b,c,d\>|\>ab=ba,bc=cb,cd=dc\rangle.$
For $i\geq 0$ set
$\alpha_{i}$ $=(ad)^{i}(aca^{-1})(ad)^{-i},$ $\beta_{i}$ $=(ad)^{i}b(ad)^{-i},$
$\mu_{i}$ $=(da)^{i}c(da)^{-i},$ $\delta_{i}$ $=(da)^{i}(dbd^{-1})(da)^{-i}.$
Then these elements are all distinct in $G$ and the subgroup $H=\langle X\rangle$ of $G$ generated by
$X=\{\alpha_{i},\;\beta_{i},\;\mu_{i},\;\delta_{i}:i\geq 0\}$
is isomorphic to the RAAG $A(P_{\infty})$ where $P_{\infty}$ is the bi-infinite line
$\alpha_{2}$$\beta_{2}$$\alpha_{1}$$\beta_{1}$$\alpha_{0}$$\beta_{0}$$\mu_{0}$$\delta_{0}$$\mu_{1}$$\delta_{1}$$\mu_{2}$$\delta_{2}$
Proof.
Since the words $\alpha_{i}$ and $\mu_{i}$ have $c$-exponent sum $1$ and the words $\beta_{i}$ and $\delta_{i}$ have $c$-exponent sum zero it follows that no word from $\{\alpha_{i},\mu_{i}:i\geq 0\}$ is equal in $G$ to any word from $\{\beta_{i},\delta_{i}:i\geq 0\}$. Note that all the words in the set $\{\alpha_{i},\mu_{i}:i\geq 0\}$ represent elements in the subgroup
$\langle a,d,c\rangle\cong\langle a\rangle\ast\langle c,d\rangle\cong\mathbb{Z}\ast(\mathbb{Z}\times\mathbb{Z})$
of $G$. Using normal forms in this free product it then follows that distinct words from the set $\{\alpha_{i},\mu_{i}:i\geq 0\}$ represent distinct elements of $G$. A similar argument working in $\langle a,b,d\rangle\cong(\mathbb{Z}\times\mathbb{Z})\ast\mathbb{Z}$ shows that distinct words from the set $\{\beta_{i},\delta_{i}:i\geq 0\}$ represent distinct elements of $G$. Hence the words in the set $X$ all represent distinct elements of the group $G$.
Clearly $\alpha_{0}=aca^{-1}$ commutes with $\beta_{0}=b$, which upon conjugation by $(ad)^{i}$ implies that $\alpha_{i}$ commutes with $\beta_{i}$ for all $i$. Also $\alpha_{i}$ commutes with $\beta_{i+1}$ since $c$ commutes with $dbd^{-1}$. Conjugating these pairs by $d$ implies that $\mu_{i}$ commutes with $\delta_{i}$, and that $\mu_{i+1}$ commutes with $\delta_{i}$. Together with the fact that $\beta_{0}=b$ commutes with $\mu_{0}=c$ this shows that words labelling adjacent vertices in $P_{\infty}$ commute in $G$.
Now let $N$ be the normal closure in $G$ of the subset $\{b,c\}\subseteq G$. It follows from [DLS91, Theorem 2] (see also the second paragraph of the proof of Theorem 3 in [DLS91]) that $N$ is isomorphic to a RAAG with generators being the set of distinct conjugates $C$ of $b$ and $c$ in $G$ by elements from the subgroup $\langle a,d\rangle\leq G$, and where the defining relators for the RAAG $N$ are given by commutators for each pair of elements from $C$ that commute with each other in $G$. It is also proved in [DLS91, Theorem 2 and Theorem 3] that the underlying graph $\Phi$ of the RAAG $N$ is a forest. By definition all of the elements from $X$ belong to the set $C$. Hence, the subgroup $H$ of $G$ generated by $X$ is in fact the subgroup of $A(\Phi)$ generated by $X$. Since $X\subseteq\Phi$, it thus follows that $H$ is isomorphic to the RAAG on the subgraph of $\Phi$ induced by $X$, i.e. $H$ is a special subgroup of $A(\Phi)$, see e.g. [Vog15, Lemma 3.2]. That is, $H$ is isomorphic to the RAAG with generating set $X$ and defining relations given by commutators for every pair of elements from $X$ that commute in $A(\Phi)$. Since $\Phi$ is a forest it follows that the only commutation relations between pairs of elements in the set $X$ are those corresponding to edges in the graph $P_{\infty}$, since any additional edge would create a cycle. This completes the proof that $H\cong A(P_{\infty})$. ∎
@@
Let $G=A(P_{4})$ be the right-angled Artin group (6). Then:
@@
Let
$G=A(P_{4})=\langle a,b,c,d\>|\>ab=ba,bc=cb,cd=dc\rangle.$
Throughout the proof we use the same notation as in the statement of Lemma 3.5. Let $F=\langle\alpha_{1},\alpha_{0},\mu_{0}\rangle\leq G$ and let $L=\langle\mu_{1},\delta_{1},\mu_{2},\delta_{2}\rangle\leq G$. It follows from Lemma 3.5 that $F$ is isomorphic to a free group of rank 3 with basis $\{\alpha_{1},\alpha_{0},\mu_{0}\}$. It also follows from Lemma 3.5 that $L$ is isomorphic to the group $A(P_{4})$, so by Theorem 3.2 the group $L$ has undecidable rational subset membership problem.
It then follows from Lemma 3.4(i) that the fixed-target submonoid membership problem for $\alpha_{0}\mu_{0}^{-1}$ in $L\ast F$ is undecidable. Indeed, if the fixed-target submonoid membership problem for $\alpha_{0}\mu_{0}^{-1}$ in $L\ast F$ were decidable then by Lemma 3.4(i) it would follow that the rational subset membership problem in $L$ is decidable, which it is not as $L\cong A(P_{4})$.
Now let
$Y=\{\alpha_{1},\alpha_{0},\mu_{0},\mu_{1},\delta_{1},\mu_{2},\delta_{2}\}\subseteq G$
and let $K=\langle Y\rangle\leq G$ be the subgroup of $G$ generated by $Y$. Then by Lemma 3.5 we have
$K$ $=$ $\langle Y\rangle=\langle\alpha_{1},\alpha_{0},\mu_{0},\mu_{1},\delta_{1},\mu_{2},\delta_{2}\rangle$
$\cong$ $\langle\alpha_{1},\alpha_{0},\mu_{0}\rangle\ast\langle\mu_{1},\delta_{1},\mu_{2},\delta_{2}\rangle\cong L\ast F.$
Since by the previous paragraph the fixed-target submonoid membership problem for $\alpha_{0}\mu_{0}^{-1}$ in $L\ast F$ is undecidable, and $K\cong L\ast F$ vie the isomorphism described above it follows that the fixed-target submonoid membership problem for $\alpha_{0}\mu_{0}^{-1}$ in $K$ is undecidable.
Note that by definition $\alpha_{0}=aca^{-1}$ and $\mu_{0}=c$ hence $\alpha_{0}\mu_{0}^{-1}=aca^{-1}c^{-1}$ in $G$. Since the fixed-target submonoid membership problem for $\alpha_{0}\mu_{0}^{-1}$ in $K$ is undecidable, and since $K$ is a finitely generated subgroup of $G$, it then follows by the contrapositive of Remark 2.1 that the fixed-target submonoid membership problem for $\alpha_{0}\mu_{0}^{-1}$ in $G$ is undecidable. But since $\alpha_{0}\mu_{0}^{-1}=aca^{-1}c^{-1}$ in $G$ this means that the fixed-target submonoid membership problem for $aca^{-1}c^{-1}$ in $G$ is undecidable. This completes the proof of part (i) of the proposition.
The proof of part (ii) of the proposition is dealt with in a similar way but by applying Lemma 3.4(ii). The details are as follows.
Maintaining the same notation as earlier in the current proof, it follows from Lemma 3.4(ii) that the semigroup intersection problem in $L\ast F$ with respect to the subsemigroup $\operatorname{Sgp}(\alpha_{0}\mu_{0}^{-1})$ is undecidable. Indeed, if the semigroup intersection problem were decidable in $L\ast F$ with respect to the subsemigroup $\operatorname{Sgp}(\alpha_{0}\mu_{0}^{-1})$ then by Lemma 3.4(ii) it would follow that the rational subset membership problem in $L$ is decidable, which it is not as $L\cong A(P_{4})$. Then since
$K=\langle Y\rangle\cong\langle\alpha_{1},\alpha_{0},\mu_{0}\rangle\ast\langle\mu_{1},\delta_{1},\mu_{2},\delta_{2}\rangle\cong L\ast F$
it follows that the semigroup intersection problem in $K$ with respect to the subsemigroup $\operatorname{Sgp}(\alpha_{0}\mu_{0}^{-1})$ is undecidable.
Since the semigroup intersection problem in $K$ with respect to the subsemigroup $\operatorname{Sgp}(\alpha_{0}\mu_{0}^{-1})$ is undecidable and since $K$ is a finitely generated subgroup of $G$, it then follows by the contrapositive of Remark 2.1 that the semigroup intersection problem in $G$ with respect to the subsemigroup $\operatorname{Sgp}(\alpha_{0}\mu_{0}^{-1})$ is undecidable. In more detail suppose, seeking a contradiction, that the semigroup intersection problem in $G$ with respect to the subsemigroup $\operatorname{Sgp}(\alpha_{0}\mu_{0}^{-1})$ is decidable. This would mean that there is an algorithm that for any finite set of words $W\subseteq\{a,b,c,d,a^{-1},b^{-1},c^{-1},d^{-1}\}^{*}$ can decide whether or not
$\operatorname{Sgp}(\alpha_{0}\mu_{0}^{-1})\cap\operatorname{Sgp}(W)=\varnothing.$
But then in particular given any finite set of words $U\subseteq(Y\cup Y^{-1})^{*}$ we can rewrite these words via the substitutions
$\alpha_{i}$ $\mapsto(ad)^{i}(aca^{-1})(ad)^{-i},$ $\beta_{i}$ $\mapsto(ad)^{i}b(ad)^{-i},$
$\mu_{i}$ $\mapsto(da)^{i}c(da)^{-i},$ $\delta_{i}$ $\mapsto(da)^{i}(dbd^{-1})(da)^{-i}$
to a finite set of words $U^{\prime}\subseteq\{a,b,c,d,a^{-1},b^{-1},c^{-1},d^{-1}\}^{*}$ representing the same set of elements of $G$ as $U$. Then by assumption we can decide whether or not
$\operatorname{Sgp}(\alpha_{0}\mu_{0}^{-1})\cap\operatorname{Sgp}(U^{\prime})=\varnothing\ \mbox{in $G$.}$
But since $\operatorname{Sgp}(U^{\prime})\leq K$ and $\operatorname{Sgp}(\alpha_{0}\mu_{0}^{-1})\leq K$ it follows that
$\operatorname{Sgp}(\alpha_{0}\mu_{0}^{-1})\cap\operatorname{Sgp}(U^{\prime})=\varnothing\ \mbox{in $G$}$
if and only if
$\operatorname{Sgp}(\alpha_{0}\mu_{0}^{-1})\cap\operatorname{Sgp}(U)=\varnothing\ \mbox{in $K$}$
where $U$ is the original finite subset $U\subseteq(Y\cup Y^{-1})^{*}$ of words over the generators of $K$. This shows that there is an algrothm that takes any finite subset $U\subseteq(Y\cup Y^{-1})^{*}$ and decides whether or not $\operatorname{Sgp}(\alpha_{0}\mu_{0}^{-1})\cap\operatorname{Sgp}(U)=\varnothing$ in $K=\langle Y\rangle$. This gives the required contradiction, since the semigroup intersection problem in $K$ with respect to the subsemigroup $\operatorname{Sgp}(\alpha_{0}\mu_{0}^{-1})$ is undecidable.
Finally, since $\alpha_{0}\mu_{0}^{-1}=aca^{-1}c^{-1}$ in $G$ this means that the semigroup intersection problem in $G$ with respect to the subsemigroup $\operatorname{Sgp}(aca^{-1}c^{-1})$ is undecidable. This completes the proof of part (ii) of the proposition. ∎
@@ 3.3. Undecidability in the braid group $\mathbf{B}_{4}$
We end this section by remarking that all the decidability results proved in this section for braid groups also hold true for a closely related object called the pure braid group $\mathbf{PB}_{n}$. Indeed, it well-known and also clear from equation (1) above (and the graphical representation of braids see [KT08, §1.2.2]) that $\mathbf{B}_{n}$ surjects onto $S_{n}$. The kernel of this homomorphism is called the pure braid group $\mathbf{PB}_{n}$, which hence has index $n!$ in $\mathbf{B}_{n}$. All of the undecidability results above for $\mathbf{B}_{n}$, were proved by means of embedding certain right-angled Artin groups $A(\Gamma)$ into $\mathbf{B}_{n}$. However, the results are equally valid for $\mathbf{PB}_{n}$, by the following reasoning. It follows from the main result of [CP01] that if $A(\Gamma)$ is a right-angled Artin group with generating set $V$ (defined by the presentation in equation (2)) then for any natural number $k\in\mathbb{N}$ with $k\geq 2$ the subgroup of $A(\Gamma)$ generated by the set of powers $\{v^{k}:v\in V\}$ is again isomorphic to the group $A(\Gamma)$. Hence by taking powers which are sufficiently large multiples of $\mathbf{B}_{n}:\mathbf{PB}_{n}]=n!$ it is not difficult to see that the following statement holds: for every finite graph $\Gamma$, there is a subgroup of $\mathbf{B}_{n}$ that is isomorphic to $A(\Gamma)$ if and only if there is a subgroup of $\mathbf{PB}_{n}$ that is isomorphic to $A(\Gamma)$. Thus, all undecidability results above for $\mathbf{B}_{n}$ with $n\geq 4$ also hold for $\mathbf{PB}_{n}$ with $n\geq 4$. To see that all the statements in Theorem 3.7 remain true for the pure braid group $\mathbf{PB}_{4}$, it simply suffices to note that $\gamma_{0}\in\mathbf{PB}_{4}$. Similarly, when $n\leq 3$, since $\mathbf{B}_{n}$ has decidable rational subset membership problem and $\mathbf{PB}_{n}\leq\mathbf{B}_{n}$ it follows that the rational subset membership problem (along with all other decision problems discussion in this section) is decidable in $\mathbf{PB}_{n}$ when $n\leq 3$.
@@ 4. Artin groups
It follows from recent work [AL21, Corollary A] that every subgroup separable Artin group lies in $A(\mathcal{S})$. The first result we will establish is the following.
Theorem 4.1.
Every group in $A(\mathcal{S})$ has decidable rational subset membership problem. In particular, every subgroup separable Artin group has decidable rational subset membership problem.
In [KSS07, Theorem 3.1] it is shown that a finitely generated group $G$ has decidable rational subset membership problem if and only if the word problem for $G$ with respect to $A$ is RID (rational intersection decidable). Here the word problem for the group is the set of all words over the finite generating set that are equal to the identity element of the group. In [LS08] Lohrey & Steinberg introduce another class called SLI (semilinear intersection) languages, and corresponding class of SLI-groups. They prove [LS08, Lemma 4] that this class is contained in the class of languages RID, so every SLI-group is an RID-group and hence has decidable rational subset membership problem.
@@
    AA has decidable submonoid membership problem;
@@
    AA has decidable rational subset membership problem;
@@
((iii) $\Rightarrow$ (ii)) By [AL24, Theorem 1.1] (which is proved using the main result of [AL21]) it follows that if $\Gamma$ does not embed any of the graphs in (a), (b), or (c) then $A(\Gamma)$ is subgroup separable which by Theorem 4.1 implies that $A$ has decidable rational subset membership problem. ((ii) $\Rightarrow$ (i)) is trivial. ((i) $\Rightarrow$ (iii)) Suppose that $\Gamma$ does embed one of the graphs of (a), (b) or (c) as an induced subgraph. Let us consider each case in turn. Case (a): In this case by [CP01] the RAAG $A(C_{4})\cong F_{2}\times F_{2}$ embeds into $A(\Gamma)$, which hence contains a fixed finitely generated submonoid (in fact subgroup) in which membership is undecidable by [Mik58]. Cases (b) and (c): In these cases it follows from the argument of the proof of [AL21, Lemma 4.2] that in each case $A(\Gamma)$ contains [^3] We note that diagram (4.1) in the proof of [AL21, Lemma 4.2] there is an edge between $x$ and $y$ that should not be there. Indeed, by the definition of $R(B)$ in that paper for there to be an edge between $x$ and $y$ one would need every vertex of $\{a,b\}$ to be connected to every vertex of $\{b,c\}$ by a $2$-labelled edge in the defining graph of the Artin group which is not true. This does not affect the validity of the proof of [AL21, Lemma 4.2], or the proof of Theorem 4.4 above, since the RAAG of the graph (4.1) with that edge removed still contains an induced subgraph isomorphic to a path of length $4$. Related to this, note that since the braid group $B_{4}\cong A(2,3,3)$ does not embed $A(C_{4})=F_{2}\times F_{2}$ it follows that (4.1) cannot contain a square as an induced subgraph. the RAAG $A(P_{4})$ and hence by [LS08] it follows that $A(\Gamma)$ contains a fixed finitely generated submonoid in which membership is undecidable. ∎
@@
Let $A=A(\Gamma)$ be an Artin group with undecidable submonoid membership problem, i.e. such that one of the graphs in (a), (b) or (c) from the statement of Theorem 4.4 embeds into $\Gamma$.
@@
    If one of the graphs in (b) or (c) embeds into $\Gamma$, then $A(P_{4})$ embeds into $A(\Gamma)$, and the following problems are all undecidable in $A(\Gamma)$: the rational subset membership problem, submonoid membership problem, fixed-target submonoid membership problem, and the subsemigroup intersection problem.
@@
    If one of the graphs in (a) embeds into $\Gamma$, then $A(C_{4})$ embeds into $A(\Gamma)$, and all the problems listed in (i) are undecidable in $A(\Gamma)$. Additionally, all the following problems are undecidable in $A(\Gamma)$: the group problem, identity problem, and subgroup membership problem.
@@
Each and every one of the following problems is decidable for an Artin group $A=A(\Gamma)$ if and only if none of the graphs in (a), (b) and (c) from the statement of Theorem 4.4 embeds into $\Gamma$: the rational subset membership problem, submonoid membership problem, fixed-target submonoid membership problem, and the subsemigroup intersection problem.
Just as for the braid group $\mathbf{B}_{4}$ we do not know exactly which Artin groups have decidable identity problem or group problem. Similarly, it is unknown exactly which Artin groups have decidable subgroup membership problem. In particular, it remains open whether the braid group $\mathbf{B}_{4}$ has decidable subgroup membership problem. Related to this, it follows from the proof of [AL21, Lemma 4.2] that $\mathbf{B}_{4}$ contains $A(C_{5})$, where $C_{5}$ is a pentagon, and it is also an open problem whether $A(C_{5})$ has decidable subgroup membership problem; see [LS08, Section 5].